When talking about aing bets, What is the diffrence between probability and odds?
Answer #1

Probability is the chance of occurring of an event expressed as a percentage or a decimal and ranges from 0 to 1 with 0 being an impossible event and 1 being a sure event.

Odds is a ratio of desired outcome to undesired outcome and ranges from 0 to infinity. There are odds in favor of an event and odds against an event.

For example, the probability of getting a head in a coin toss is 1/2 or 50% while the odds in favor a head is 1/1 =1:1. In a dice throw, the probability of getting 2 is 1/6 while the odds in favor of 2 is 1/5 =1:5 and odds against 2 is 5/1 =5:1

The probability of an event A is:

P(A) =Favorable outcomes/Total outcomes

The Odds in favor of A is:

Odds =Favorable outcomes/Unfavorable outcomes